OpenStudy (anonymous):

2, 10, 30, 68, 130, _____

OpenStudy (lgbasallote):

^?

OpenStudy (anonymous):

answer's 222 i donno how

OpenStudy (anonymous):

ok ok got it thanks

OpenStudy (anonymous):

Find the difference between numbers like: 10-2=8 30-10=20 68-30=38 130-68=62 again find the diff. 20-8=12 38-20=18 62-38=24 the diff. is increasing by 6. so the next diff. should be 30. add this 30 to find the first diff. which is 62+30=92. add this 92 to 130 you will get 222.

OpenStudy (anonymous):

u got it sweetie